OpenStudy (anonymous):

Please help! a+b=5 b+c=-2 c+a=7 Solve

jimthompson5910 (jim_thompson5910):

you can isolate 'a' in the third equation c+a=7 c+a-c=7-c ... subtract 'c' from both sides a = 7-c then plug this into the first equation a+b=5 7-c+b=5 ... replace 'a' with 7-c 7+b-c = 5 7+b-c-7 = 5-7 ... subtract 7 from both sides b-c = -2

jimthompson5910 (jim_thompson5910):

after doing all that, we end up with this system of equations b+c=-2 b-c = -2 do you see what to do next?

OpenStudy (anonymous):

I understand so far but I don't know what to do next.

OpenStudy (texaschic101):

try adding the 2 equations and what do you get ?

OpenStudy (texaschic101):

notice that the c's cancel

OpenStudy (anonymous):

b=-4 ?

OpenStudy (texaschic101):

yep

jimthompson5910 (jim_thompson5910):

you forgot to add the b terms b+b = 2b

OpenStudy (texaschic101):

oops..my bad....sorry, I missed that

jimthompson5910 (jim_thompson5910):

so if 2b = -4, then b = ???

OpenStudy (anonymous):

b=-2

OpenStudy (texaschic101):

yes....now sub -2 in for b in the other equations to find the other letters

OpenStudy (anonymous):

a=7 and c=0 ?

jimthompson5910 (jim_thompson5910):

so we have this claim a = 7 b = -2 c = 0 let's test that claim with each equation Testing equation 1: a+b = 5 7+(-2) = 5 ... plug in a = 7 and b = -2 5 = 5 ... true Testing equation 2 b+c = -2 -2+0 = -2 ... plug in b = -2 and c = 0 -2 = -2 ... true Testing equation 3 a+c = 7 7+0 = 7 ... plug in a = 7 and c = 0 7 = 7 ... true All three equations are true when a = 7, b = -2 and c = 0. So the solution (a,b,c) = (7,-2,0) is confirmed

jimthompson5910 (jim_thompson5910):

whenever you get a possible solution, it's always a good idea to check that possible solution back into the original equation(s) to see if it's a true solution or not

OpenStudy (anonymous):

okay thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome